OpenStudy (anonymous):

Determine dy/dx for the following: 1.y=e^3x.cos(x^2+1) 2.y=ln(sqrt e^3x(x^2-1))

OpenStudy (dape):

Just remember the chain and product rule, for the first problem we have \[ \frac{dy}{dx} =\frac{d}{dx}[e^{3x}\cos(x^2+1)] \\ =\frac{d}{dx}[e^{3x}]\cos(x^2+1)+e^{3x}\frac{d}{dx}[\cos(x^2+1)] \\ =3e^{3x}\cos(x^2+1)+e^{3x}(-\sin(x^2+1)*2x) \\ =e^{3x}(3\cos(x^2+1)-2x\sin(x^2+1)) \]

OpenStudy (dape):

Try to do the second problem yourself

OpenStudy (anonymous):

Ok thank you,your a star! Just wanted to make sure my answers are right :)

OpenStudy (dape):

No problem, glad to help!